If a consumers objective function is given by U=X1aX2subject to P1X1 + P2X2 = m. Establish the levels of X1 and X2 that maximizes the consumers satisfaction.


1
Expert's answer
2021-06-01T13:49:45-0400

Solution


U=X1aX2b

M=P1X1+P2X2

MUx1="\frac{DU}{DX1}=aX"1a-1X2b


MUX2="\frac{DU}{DX_2}=bX_1^aX^{b-1}"

maximizing utility


"\\frac{MU x1}{MU x2\ufeff}" ="\frac{P 1}{P2}"


"\frac{aX_1^{a-1}x_2^b}{bX_1^aX_2^{b-1}}" ="\frac{P_1}{P_2}"


solve for X2

"\frac{aX_2^bX_2^{-b+1}}{ bX_1^aX_1^{-a+1}}" ="\frac{P_1}{P_2}"


"\frac{aX_2}{bX_1}" ="\frac{P_1}{P_2}" X2= "\frac{bP_1}{aP_2}"


M=P1X1+P2X2

substituting X2


M= P1X1+P"\frac{bP_1}{aP_2}" X1


M="\frac{aP_1X_1}{a}" +"\frac{bP_1X_1}{a}"


M="\frac{(a+b)P_1}{a}" X1


X1"\frac{aM}{(a+b)P_1}"


X2"\frac{bP_1}{aP2}" x "\frac{aM}{(a+b)P_1}"


X2="\frac{bM}{(a+b)P_2}"